I'm problems with the following problem in understanding it.


Find a polynomial with leading coefficient 1 and degree 3 that has -1, 1, and 3 as roots.


It seems pretty easy and I think the answer is x^3-3x^2-x-3, but that just from looking at it. I'm not sure how to actually do the problem.


Thanks for any help

Answer by xcentaur(357)


Now let the polynomial be 

x^3-ax^2+bx-c=0

Let the roots be p,q,and r.

Then a=p+q+r
     b=pq+qr+pr
     c=pqr

since the roots are given as -1,1 and 3,we get

a=-1+1+3=3

b=(-1*1)+(1*3)+(-1*3)=-1

c=(-1)(1)(3)=-3

substituting,
x^3-ax^2+bx-c=0

x^3-(3)x^2+(-1)x-(-3)=0

x^3-3x^2-x+3=0

Now this can be verified for a polynomial as: (x-p)(x-q)(x-r)=f(x)
(x-(-1))(x-(+1))(x-(+3)) (x-1)(x+1)(x-3) On expanding we get the expression that we just worked out. Thus we know that our polynomial is indeed correct.


Hope this helps,
Prabhat